# Super-Resolution with Structured Motion

### v1

@@ Abstract
We consider the limits of super-resolution using imaging constraints. Due to various theoretical and practical limitations, reconstruction-based methods have been largely restricted to small increases in resolution. In addition, motion-blur is usually seen as a nuisance that impedes super-resolution. We show that by using high-precision motion information, sparse image priors, and convex optimization, it is possible to increase resolution by large factors. A key operation in super-resolution is deconvolution with a box. In general, convolution with a box is not invertible. However, we obtain perfect reconstructions of sparse signals using convex optimization. We also show that motion blur can be helpful for super-resolution. We demonstrate that using pseudo-random motion it is possible to reconstruct a high-resolution target using a single low-resolution image. We present numerical experiments with simulated data and results with real data captured by a camera mounted on a computer controlled stage.
@@ 1 Introduction
One of our motivations is to implement a compressed sensing system for super-resolution. Using a high-resolution CCD/CMOS sensor, it may be possible to reconstruct an extremely high-resolution image using an appropriate optical and computational imaging system.
Classical reconstruction-based methods relate a collection of low-resolution images to a high-resolution image using a model of the imaging process. Such methods have been largely restricted to small increases in resolution due to theoretical and practical limitations [1, 2]. In addition, motion blur is usually seen as a nuisance that impedes super-resolution [3].
We show that by using multiple static images acquired in a grid of locations and sparse image priors, it is possible to increase resolution by large factors. We note a connection between super-resolution and deconvolution with a box filter, which helps to understand the ambiguities in the super-resolution problem. In particular, the low-resolution images determine most of the Fourier components of the super-resolution image and, in this setting, sparse images can be recovered using convex optimization.
We also show that by using pseudorandom motion it may be possible to reconstruct a high-resolution target using a single low-resolution image. Figure 1 illustrates this idea. In this case, we capture an image while the camera/sensor undergoes a large vibration. By solving a convex optimization problem with the imaging constraints and sparsity prior, we obtain an essentially perfect reconstruction of the high-resolution target from a single low-resolution image.
@@
Super-resolution imaging systems have many potential applications. In some settings, there are practical and physical limitations that lead to low-resolution sensors. For the case of visible light photography with high-resolution sensors, our methods could be used to produce gigapixel images using a small camera. High-resolution imaging has many applications in science and in archival/preservation work (including for art, documents, and biological samples). Furthermore, there are many applications where a camera might be constantly moving, such as in aerial imaging.
We implemented a system to demonstrate some of our ideas using a commonly available camera and computer controlled motion stage. The experiments in Section 6 demonstrate results using both static images and images taken with the camera moving at constant velocity.
Throughout the paper, we focus on cases where the effective resolution is limited by the sensor and not by the optical path (diffraction). Even with high-resolution sensors, super-resolution can be useful for obtaining a wider field of view with an appropriate lens. The interlacing operation described in Section 6.2 is reminiscent of a panorama, but interlaces a set of pictures taken with small motion instead of tiling pictures taken with large motion, which leads to a more compact design for a high-resolution camera. The Wiener filter approach for reconstruction is very efficient and yields a significant improvement in resolution in our experiments. In practice, large images can also be processed in parallel via tiling.
We assume throughout the paper that the sensor motion is known, either because it is carefully planned and executed by a high-accuracy actuator or because we have a high-accuracy readout of the sensor position while it is being moved through some other means (such as vibration). For the experiments in Section 6.3, we estimate the camera trajectory during an exposure using a landmark.
@@ 2 Related Work
The super-resolution problem has a long history (see, e.g.[4, 5, 1, 2]). From a theoretical point of view, the results in [1] and [2] are largely negative, showing that ambiguity in the super-resolution problem grows quickly with the desired increase in resolution. In contrast, here we note that with a particular sampling scheme this ambiguity is concentrated on relatively few Fourier coefficients. Previous work in compressed sensing [6, 7] has shown that sparse image priors can be used to resolve such ambiguities.
@@
Recently, some imaging schemes have been proposed, not to achieve super-resolution, but to remove motion-related artifacts. For example, parabolic lens motion [13], induced by constant acceleration, generates a PSF that is invariant to object velocity. Additionally, coded blur generated by fluttering a camera shutter [14, 15] has been shown to help in removing motion blur. This leads to a surprising idea that it may be beneficial to extend an exposure with a coded shutter to compensate for motion blur.
@@
In this paper, we focus on the use of total variation for reconstruction. In practice, alternative image priors, such as ones based on self-similarity or deep networks, could also be implemented using Plug-and-Play methods [19].
@@ 3 Imaging Model
We will also consider the case where the sensor moves while the image is captured. If the sensor follows a trajectory $(x(t),y(t))$, we capture an image $I$ with
@@
Note that translating the sensor by $z$ is equivalent to translating $g$ by $-z$. Let $q(x,y)$ denote the occupancy map defined by the trajectory $(-x(t),-y(t))$. The image $I$ captured by the moving sensor is equal the image of $g\otimes q$ captured by a static sensor.
@@ 4 Grid of Images
with $k$ and $l$ integers. In this case, the low-resolution pixels in $I$ are aligned with the high-resolution pixels in $J$. In particular, each value in $I$ is the sum of $f^{2}$ values in $J$,
@@
Suppose we capture $f^{2}$ different low-resolution images $\{I_{k,l}\}$ by translating the sensor over an $f\times f$ grid of high-resolution locations,
@@
Capturing such images involves moving the sensor to sub-pixel locations defined with a step size $\Delta/f$. The total vertical and horizontal displacement required to capture all images is less than $\Delta$.
@@ 4.1 Interlacing
The $f^{2}$ different images $\{I_{k,l}\}$ can be rearranged into a single $fn\times fm$ image $H$ such that $H=J\otimes B$, where $B$ is a box of width $f$. Since each image $I_{k,l}$ is a decimation of $J\otimes B$, the rearrangement interlaces the captured images:
@@
Figure 2 illustrates the process where we capture multiple low-resolution images and interlace the result to obtain a single “blurry” image $H$. The interlaced image $H$ has $fn\times fm$ pixels and defines the imaging constraints on $J$ through the relation $H=J\otimes B$.
@@ 4.2 Deconvolution with a Box
Note that a box filter is not a good smoothing or low-pass filter. Convolution with a box wields a signal that is visually blurry but still has quite a lot of high-frequency information. In Figure 3(a) we compare the Fourier transform of a one-dimensional box and a Gaussian of similar width. The Fourier transform of a Gaussian is a Gaussian and the Fourier transform of a box is a sinc. In particular, the magnitude of the Fourier transform of the box decays much slower than the magnitude of the Fourier transform of a Gaussian.
For the case of a two-dimensional discrete (periodic) signal, we illustrate the magnitude of the DFT of $B$ when $f=4$ in Figure 3(b). Note that there are relatively few coefficients in the DFT of $B$ with magnitude close to 0. We can therefore directly recover most of the coefficients in the Fourier transform of $J$ from the Fourier transform of $H$, except for isolated frequencies where the Fourier transform of $B$ is zero or close to zero.
@@
Previous work [1, 2] showed that the ambiguity in $J$ grows exponentially with $f$. We note, however, that even for large factors $f$, the constraints defined by $J=H\otimes B$ identify most of the Fourier coefficients of $J$.
The number of times the Fourier transform of a discrete periodic box of width $f$ crosses zero is $O(f)$. Therefore, the ambiguity in $J$ is concentrated on relatively few Fourier coefficients. In this setting, sparse image priors and convex optimization are known to yield surprisingly good reconstructions [6, 7].
For the specific case of deconvolution with a box, we have shown [22] that convex optimization using $\ell_{1}$ regularization can lead to perfect reconstruction of sparse one-dimensional signals.
@@ 4.3 Wiener Deconvolution
Using a Gaussian prior on the Fourier coefficients of $J$ leads to a fast Wiener deconvolution algorithm via the FFT and an inverse transform. The result corresponds to a MAP estimate of the super-resolution image $J$. The prior is defined by a single parameter $\gamma$ used to shrink the estimated Fourier coefficients.
@@
4.4 Sparse Reconstruction
As mentioned above, previous work has shown that sparse signals can often be recovered from incomplete frequency information [6, 7].
In practice, we regularize the reconstruction of the super-resolution image $J$ using the total variation (TV) measure, which leads to a convex optimization problem,
@@ 5 Super-resolution by Deblurring
First, we note that a small amount of motion blur can be helpful for super-resolution. Consider the problem of localizing a point source in one dimension; without motion, we can only localize the point source to the pixel size. Now suppose we move the sensor/camera with constant velocity by one pixel while taking a picture. Figure 5 illustrates this situation. The point source is spread over two pixels and we observe two non-zero values $I[k]$ and $I[k+1]$. Together, these measurements determine both the total light intensity and the precise point source location,
@@
Figure 5: We can localize a point source to sub-pixel location if the camera undergoes a motion of one pixel during capture. A motion of one pixel spreads the source to two pixels, and the relative values determine the source location. On the left, we have the continuous brightness function and the image captured with a static sensor. On the right, the sensor moves with constant velocity by one pixel during capture.
We also note that for “sparse” scenes, we can use large motions to (in essence) take multiple low-resolution images in different areas of the sensor. Shown below in Figure 6 is a simulation where we vibrate a sensor while capturing an image, which leads to the somewhat counterintuitive notion that shaking the camera during an exposure can lead to sharper pictures.
@@
As in the previous section, let $J$ be a super-resolution image generated by a virtual sensor at the origin. Now we let $I$ be an image captured by a moving sensor. Let $(k(t),l(t))$ for $0\leq t\leq 1$ denote the trajectory of the sensor over the grid of high-resolution pixel locations. That is, at time $t$, the sensor is at location $(k(t)(\Delta/f),l(t)(\Delta/f))$.
Let $Q$ be the occupancy map of $(-k(t),-l(t))$. That is, $Q[k,l]$ is the total fraction of time this trajectory spends at location $(k,l)$. When the sensor undergoes a continuous trajectory we use linear interpolation to define a discrete occupancy map $Q$. Let $B$ be a box of width $f$. The relationship between $J$ and $I$ is captured by a decimation by factor $f$ of the sequential convolution of $J$ with $Q$ and $B$. The convolution of $J$ with $Q$ accounts for the sensor motion, while the convolution with $B$ and decimation accounts for the relative size of the physical and virtual pixels.
@@
This system of equations has many fewer constraints than variables, but it can lead to perfect reconstructions with an appropriate image prior as we demonstrate in simulations below.
In some applications we may have a sequence of low-resolution images $I_{1},\ldots,I_{s}$ taken over one or multiple trajectories. Let $O_{j}$ denote the occupancy map of the trajectory followed during the acquisition of $I_{j}$.
We can reconstruct a single super-resolution image using the data from multiple trajectories and a sparsity prior. To estimate a sparse image image use $\ell_{1}$ regularization,
@@
Figure 6: (a) High-resolution targets with 10 random characters each. (b) Images recorded by a static low-resolution sensor. (c) Sensor trajectory (vibration on the top and random translations/shifts on the bottom). (d) Low-resolution images recorded with a sensor undergoing large motions. (e) Reconstruction of the high-resolution target from the blurry images in (d).
@@ 5.1 Simulations
We evaluate the quality of the reconstructions obtained using a single exposure when the sensor moves in different trajectories.
Figure 6(a) shows two examples of the high-resolution targets we used for the experiments. Each target is a $128\times 128$ image with several random characters placed in random locations. Figure 6(b) shows the results of taking a low-resolution picture of each target using a $32\times 32$ sensor. In this case, there is no sensor motion. We can see that with such a low-resolution picture, the characters become completely unreadable.
Figure 6(c) shows the results of imaging the two targets from Figure 6(a) with a moving sensor. In each case, the sensor undergoes a large motion. Figure 6(d) shows the reconstructions of the targets from a single low-resolution image, captured while the sensor is moving. In these examples, we obtain a perfect super-resolution image from a single low-resolution image by leveraging motion blur.
For the experiments in this section, we treat the high-resolution target as periodic. The sensor motion therefore leads to a cyclic convolution between the target and the sensor occupancy map. We used the $\ell_{1}$ norm of the target image $J$ to regularize the reconstructions.
@@
    1000 random displacements (shifts). In this case, the trajectory is a sequence of 1000 random locations. Each location is selected independently from a uniform distribution in the high-resolution grid.
@@
    Random walk with 1000 steps (walk). In this case, the sensor undergoes a random walk in the grid of high-resolution locations.
@@
    Vibration (vibration). The $x$ and $y$ coordinates of the sensor are defined by two sinusoidals.

### v2

@@ Abstract
We consider the limits of super-resolution using imaging constraints. Due to various theoretical and practical limitations, reconstruction-based methods have been largely restricted to small increases in resolution. In addition, motion-blur is usually seen as a nuisance that impedes super-resolution. We show that by using high-precision motion information, sparse image priors, and convex optimization, it is possible to increase resolution by large factors. A key operation in super-resolution is deconvolution with a box. In general, convolution with a box is not invertible. However, we obtain perfect reconstructions of sparse signals using convex optimization. We also show that motion blur can be helpful for super-resolution. We demonstrate that using pseudo-random motion it is possible to reconstruct a high-resolution target using a single low-resolution image. We present numerical experiments with simulated data and results with real data captured by a camera mounted on a computer-controlled stage.
@@ 1 Introduction
One of our motivations is to implement a compressed sensing system for super-resolution. Using a high-resolution CCD/CMOS sensor, it may be possible to reconstruct an extremely high-resolution image using an appropriate optical and computational system.
Classical methods for super-resolution relate a collection of low-resolution images to a high-resolution image using a model of the imaging process. Such methods have been largely restricted to relatively small increases in resolution due to theoretical and practical limitations [1, 2]. In addition, motion blur is usually seen as a nuisance that impedes super-resolution [3].
Despite the theoretical limits previously described in the literature, we show that it is possible to increase resolution by large factors using sparse image priors. We note a connection between super-resolution and deconvolution with a box filter, which helps to understand the ambiguities in the super-resolution problem. In particular, a collection of low-resolution images determines most of the Fourier components of the super-resolution image and, in this setting, sparse images can be recovered using convex optimization.
We also show that by using motion blur, it is possible to reconstruct a high-resolution target from a single low-resolution image. By moving the camera or sensor during an exposure, we capture a superposition of multiple low-resolution images. Figure 1 illustrates the idea. In this example where we capture an image while the camera/sensor undergoes a large vibration. We record a single low-resolution image and reconstruct a high-resolution target by solving a convex optimization problem.
@@
Super-resolution imaging systems have many potential applications. In some settings, there are practical and physical limitations that lead to low-resolution sensors. Furthermore, there are many applications where a camera might be constantly moving, such as in aerial imaging. High-resolution imaging has many applications in science and in archival/preservation work (including for art, documents, and biological samples). For example, [4] describes a large format tile-scan camera to image static scenes for cultural-heritage preservation.
For the case of visible light photography with high-resolution sensors, our methods could be used to produce gigapixel images using a small camera. Pixel shift cameras are commercially available and already use sensor motion to overcome the resolution limit induced by color filters. The same technology could be used to implement the imaging methods described here.
We implemented a system to demonstrate some of our ideas using a commonly available camera and computer-controlled motion stage. The experiments in Section 6 demonstrate results using both static images and images taken with the camera moving at a constant velocity.
Throughout the paper, we focus on cases where the effective resolution is limited by the sensor and not by the optical path (diffraction). Even with high-resolution sensors, super-resolution can be useful for obtaining a wider field of view with an appropriate lens. The interlacing operation described in Section 6.2 is reminiscent of a panorama, but interlaces a set of pictures taken with small motion instead of tiling pictures taken with large motion, which leads to a more compact design for a high-resolution camera.
We assume throughout the paper that the sensor motion is known, either because it is carefully planned and executed by a high-accuracy actuator or because we have a high-accuracy readout of the sensor position while it is being moved through some other means (such as vibration). In practice, the motion information should be accurate to the target pixel resolution. In Section 6.3, we use a landmark (sharp edge) to estimate a high-resolution camera position directly from a low-resolution image.
@@ 2 Related Work
The super-resolution problem has a long history (see, e.g.[5, 6, 1, 2]). From a theoretical point of view, the results in [1] and [2] are largely negative, showing that ambiguity in the super-resolution problem grows quickly with the desired increase in resolution. In contrast, here we note that with a particular sampling scheme, this ambiguity is concentrated on relatively few Fourier coefficients. Previous work in compressed sensing [7, 8] has shown that sparse image priors can be used to resolve such ambiguities.
@@
Recently, some unusual imaging schemes have been proposed, not to achieve super-resolution, but to remove motion-related artifacts. For example, parabolic lens motion [14], induced by constant acceleration, generates a PSF that is invariant to object velocity. Additionally, fluttering a camera shutter [15, 16] has been shown to help in removing motion blur. This leads to a surprising idea that it may be beneficial to extend an exposure with a coded shutter to compensate for motion blur.
@@
In this paper, we focus on the use of total variation for reconstruction. In practice, alternative image priors, such as ones based on self-similarity or deep networks, could also be implemented using Plug-and-Play methods [20]. Total variation was also used as prior for super-resolution in [6].
@@ 3 Imaging Model
We will also consider the case where the sensor moves while the image is captured. If the sensor follows a trajectory $p(t)=(x_{0}(t),y_{0}(t))$, we capture an image $I$ with
@@
Note that translating the sensor by $z$ is equivalent to translating $g$ by $-z$. Let $q(x,y)$ denote the occupancy map defined by the trajectory $-p(t)$. That is, $q(x,y)$ is the fraction of time the trajectory spends at position $(x,y)$. The image $I$ captured by the moving sensor is equivalent to the image of $g\otimes q$ captured by a static sensor.
@@ 4 Grid of Images
with $k$ and $l$ integers. In this case, the low-resolution pixels in $I$ are aligned with the high-resolution pixels in $J$. Therefore, each value in $I$ is the sum of $f^{2}$ values in $J$,
@@
Now suppose we capture $f^{2}$ images $\{I_{k,l}\}$ by translating the sensor over an $f\times f$ grid of high-resolution locations,
@@
Capturing such images involves moving the sensor to sub-pixel locations defined by a step size $\Delta/f$. The total vertical and horizontal displacement required to capture all images is less than $\Delta$.
@@ 4.1 Interlacing
The images $\{I_{k,l}\}$ can be rearranged into a single $fn\times fm$ image $H$ such that $H=J\otimes B$, where $B$ is a box of width $f$. Since each image $I_{k,l}$ is a decimation of $J\otimes B$, the rearrangement interlaces the captured images:
@@
Figure 2 illustrates the process where we capture multiple low-resolution images at sub-pixel shifts and interlace the result to obtain a single image $H$. The interlaced image $H$ has $fn\times fm$ pixels and defines the imaging constraints on $J$ through the relation $H=J\otimes B$.
@@ 4.2 Deconvolution with a Box
Note that a box filter is not a good smoothing or low-pass filter. Convolution with a box yields a signal that is visually blurry but still has quite a lot of high-frequency information. In Figure 3(a) we compare the Fourier transform of a one-dimensional box and a Gaussian of similar width. The Fourier transform of a box is a sinc, and the Fourier transform of a Gaussian is a Gaussian. Although the sinc has isolated zeros, the magnitude of the Fourier transform of the box decays much slower than the magnitude of the Fourier transform of a Gaussian.
For a two-dimensional discrete (and periodic) signal, we illustrate the magnitude of the DFT of a box of width $4$ in Figure 3(b). Note that there are relatively few coefficients in the DFT of $B$ with magnitude close to 0. We can therefore directly recover most of the coefficients in the Fourier transform of $J$ from the Fourier transform of $H$, except for isolated frequencies where the Fourier transform of $B$ is zero or close to zero.
@@
This is exactly the setting in which sparse image priors and convex optimization are known to yield surprisingly good (“perfect”) reconstructions [7, 8].
For one-dimensional sparse signals, deconvolution with a box using $\ell_{1}$ regularization leads to perfect reconstructions [23]. This is true despite the ambiguity in the imaging constraints due to the unknown Fourier coefficients. The key here is the choice of prior which selects a particular solution among all solutions that agree (or nearly agrees in the case of noisy data) with the measurements.
@@
4.3 The Limits of Super-Resolution
Previous work [1] considered the setting discussed here and showed that the ambiguity in the super-resolution problem grows quickly with the super-resolution factor $f$. This was characterized in terms of the volume of the set of solutions that are compatible with a set of images and the condition number of a resulting linear system.
We note, however, that the negative results in [1] do not account for the use of a prior to select a particular solution among the ones that are compatible (or nearly compatible) with the measured data. For the experiments with imaging constraints in [1] the ambiguity was resolved with a quadratic smoothness prior, while we use total variation.
The difference between using a quadratic smoothness prior and total variation may seem subtle at first degree, but one of the most striking results in compressed sensing is that priors that induce sparsity, like total variation, can lead to perfect reconstructions with under-determined linear systems [7]. The fact that natural images have the appropriate sparsity structure is a surprising phenomenon of its own. Total variation is also known to preserve/recover sharp edges in image restoration (see, e.g. [24]).
Figure 5 compares a result that minimizes a quadratic smoothness prior (as in [1]) to a result that minimizes the total variation. In this case there is no measurement error and both solutions satisfy the imaging constraints defined by $H=J\otimes B$. This illustrates how the solution is ambiguous (as expected) and how the choice of image prior can have a significant impact on the quality of the results. Moreover, the example illustrates that a low-level prior is sufficient to resolve high-resolution details.
(a) Interlaced images ($H$)
(b) Quadratic prior
(c) TV prior
Figure 5: Deconvolution of $H$ with a quadratic smoothness prior leads to an overly smooth image and ringing artifacts while using a TV prior yields an almost perfect reconstruction.
@@ 4.4 Wiener Deconvolution
Using a Gaussian prior for the Fourier coefficients of the high-resolution image leads to a very fast Wiener deconvolution algorithm via the FFT and an inverse transform.
The Wiener filter computes a MAP estimate of the super-resolution image $J$. The algorithm is defined by a single parameter $\gamma$ used to shrink the estimated Fourier coefficients.
@@
4.5 Sparse Reconstruction with TV prior
As discussed above, previous work has shown that sparse signals can often be perfectly or nearly perfectly recovered from incomplete frequency information [7, 8].
In practice, we regularize the reconstruction of the super-resolution image $J$ using the total variation (TV) measure. This leads to a convex optimization problem,
@@ 5 Super-resolution by Deblurring
First, we note that a small amount of motion blur can be helpful for super-resolution. Consider the problem of localizing a point source in one dimension; without motion, we can only localize the point source to the pixel size. Now suppose we move the sensor/camera with constant velocity by exactly one pixel while taking a picture. Figure 6 illustrates this situation. The motion leads to a convolution of the brightness function with a box. The point source is spread over two pixels and we observe two non-zero values $b=I[k]$ and $c=I[k+1]$. Together, these measurements determine both the total light intensity and the sub-pixel point source location,
@@
Figure 6: We can localize a point source to sub-pixel location if the camera undergoes constant velocity motion of one pixel during capture. The motion spreads the source to two pixels, and the relative values determine the source location. On the left, we have the continuous brightness function and the image captured with a static sensor. On the right, the sensor moves with constant velocity by one pixel during capture. The effect of the sensor motion is equivalent to a convolution of the brightness function with a box.
We also note that for “sparse” scenes, we can use large motions to (in essence) take multiple low-resolution images in different areas of the sensor. Figure 7 illustrates a simulation where the sensor undergoes a large magnitude vibration (top row) while capturing an image. This leads to the somewhat counterintuitive notion that shaking the camera during an exposure can lead to sharper pictures.
@@
As in the previous section, let $J$ be a super-resolution image generated by a virtual sensor at the origin. Now let $I$ be an image captured by a moving sensor. Let $p(t)=(k(t),l(t))$ for $0\leq t\leq 1$ denote a sensor trajectory on the grid of high-resolution pixel locations.
Let $B$ be a two-dimensional box of width $f$ and $Q$ be the occupancy map of the trajectory $-p(t)$. When the sensor undergoes continuous motion we use linear interpolation to define a discrete high-resolution occupancy map $Q$. The relationship between $J$ and $I$ is captured by a decimation by factor $f$ of the sequential convolution of $J$ with $Q$ and $B$. The convolution of $J$ with $Q$ accounts for the sensor motion, while the convolution with $B$ and decimation accounts for the relative size of the physical and virtual pixels.
@@
Due to the decimation by $f$, this system of equations has many fewer constraints than variables. The convolution with $B$ also introduces ambiguities. Nonetheless, these constraints can lead to near perfect reconstructions with an appropriate choice of image prior as we demonstrate in simulations below.
In some applications we may have a sequence of low-resolution images $I_{1},\ldots,I_{s}$ taken over one or multiple trajectories. Let $Q_{j}$ denote the occupancy map of the trajectory followed during the acquisition of $I_{j}$.
We can reconstruct a high-resolution image using the data from multiple motion blurred images simultaneously. To estimate a sparse image using $\ell_{1}$ regularization we solve the convex optimization problem,
@@
Figure 7: (a) High-resolution targets with 10 random characters. (b) Images recorded by a static low-resolution sensor. (c) Sensor trajectory (vibration on the top and random translations on the bottom). (d) Low-resolution images recorded by a moving sensor. (e) Reconstruction of the high-resolution target from the image recorded by the moving sensor.
@@ 5.1 Simulations
To demonstrate the use of motion blur for super-resolution, we simulated images obtained using different sensor trajectories. For the experiments in this section, we treat the high-resolution target as periodic. The sensor motion therefore leads to a cyclic convolution between the target and the sensor occupancy map. We used the $\ell_{1}$ norm of the target image $J$ to regularize the reconstructions.
Figure 7(a) shows two examples of the high-resolution targets used in the experiments. Each target is a $128\times 128$ image with several random characters placed in random locations. Figure 7(b) shows the results of taking a low-resolution picture of each target using a $32\times 32$ sensor. In this case, there is no sensor motion. We can see that with such a low-resolution picture, the characters become completely unreadable.
Figure 7(d) shows the results of imaging the two targets from Figure 7(a) with a moving sensor. In each case, the sensor undergoes a large motion. Figure 7(e) shows the reconstructions of the targets from the low-resolution image captured while the sensor is moving. In these examples, we obtain a nearly perfect high-resolution image from a single low-resolution image by leveraging motion blur.
@@
    Vibration (vibration). The $x$ and $y$ coordinates of the moving sensor are defined by two sinusoidals with large magnitude. (Figure 7 top row)
@@
    1000 random displacements (shifts). In this case, the trajectory is a sequence of 1000 random locations. Each location is selected independently from a uniform distribution in the high-resolution grid. (Figure 7 bottom row)
@@
    Random walk with 1000 steps (walk). In this case, the sensor undergoes a random walk in the grid of high-resolution locations. (Not shown)